# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 9.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.184184184The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 9.5 Discussion
Stochastic-game Q-learning and CFR target complementary game classes: simultaneous-move games with observable payoffs and extensive-form games with private information, respectively. The simulations confirm convergence to known equilibria in both settings without encoding economic structure into the algorithms.
@@ 10 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where, unlike the simulator-based training of preceding chapters, the agent learns directly from interactions with real customers and exploration is a real cost that must be balanced against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.186186186Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 10.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.187187187The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm188188188Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 10.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.189189189The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.190190190This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 10.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.191191191Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 10.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.192192192If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 10.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 10.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.193193193Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.194194194The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.195195195For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 10.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.196196196The regret benchmark here differs from Section 10.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).197197197EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).198198198The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When a modeler specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.199199199Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.

### 24 Mar 2026 v5

@@ 9.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.186186186The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 9.5 Discussion
Stochastic-game Q-learning and CFR target complementary game classes: simultaneous-move games with observable payoffs and extensive-form games with private information, respectively. The simulations confirm convergence to known equilibria in both settings without encoding domain structure into the algorithms.
@@ 10 Bandits and Dynamic Pricing
This chapter focuses on in-field reinforcement learning (Section 2), where, unlike the simulator-based training of preceding chapters, the agent learns directly from interactions with real customers and exploration is a real cost that must be balanced against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.188188188Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 10.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.189189189The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm190190190Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 10.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.191191191The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.192192192This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 10.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.193193193Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 10.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.194194194If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 10.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 10.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.195195195Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.196196196The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to an in-field deployment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.197197197For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 10.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.198198198The regret benchmark here differs from Section 10.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024a) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).199199199EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).200200200The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When a modeler specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.201201201Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.